# Aryaman Arora  Neil Rathi  Nikil Roashan Selvam  Róbert Csordás Dan JurafskyChristopher Potts

### v2

@@ Abstract
State space models (SSMs) for language modelling promise an efficient and performant alternative to quadratic-attention Transformers, yet show variable performance on recalling basic information from the context. While performance on synthetic tasks like Associative Recall (AR) can point to this deficiency, behavioural metrics provide little information as to why—on a mechanistic level—certain architectures fail and others succeed. To address this, we conduct experiments on AR and find that only Transformers and Based SSM models fully succeed at AR, with Mamba a close third, whereas the other SSMs (H3, Hyena) fail. We then use causal interventions to explain why. We find that Transformers and Based learn to store key–value associations in-context using induction heads. By contrast, the SSMs compute these associations only at the last state, with only Mamba succeeding because of its short convolution component. To extend and deepen these findings, we introduce Associative Treecall (ATR), a synthetic task similar to AR based on PCFG induction. ATR introduces language-like hierarchical structure into the AR setting. We find that all architectures learn the same mechanism as they did for AR, and the same three models succeed at the task. These results reveal that architectures with similar accuracy may still have substantive differences, motivating the adoption of mechanistic evaluations.
@@ 1 Introduction
Transformers with quadratic attention remain the dominant architecture in language modelling despite numerous proposed efficient alternatives. Most notably, state space models (SSMs) achieve impressive perplexities and benchmark scores [e.g. Gu and Dao, 2024]. Yet, SSMs exhibit deficiencies that benchmarks often fail to capture; for example, they struggle to perform retrieval, i.e. copying from the context Jelassi et al. [2024], Wen et al. [2024], Waleffe et al. [2024], Bick et al. [2025].
@@
Controlled synthetic tasks can make these limitations clear by isolating specific capabilities and enabling expressive experimentation at small scales across architectures. Particularly, much work has used the associative recall (AR) task as a testbed for studying in-context retrieval across architectures. In turn, AR has informed the design of novel LM architectures [e.g. Based; Arora et al., 2024b].
@@
Yet performance on synthetic tasks is measured solely via behavioural metrics like task accuracy. This is a missed opportunity: an advantage of these synthetic tasks is that they are designed to isolate a specific behaviour that implicates a mechanistic solution. For example, language models should solve AR by storing key–value associations in-context at the value, a mechanism termed the induction head in Transformers Olsson et al. [2022], Fu et al. [2023]. We should therefore directly check whether each architecture learns induction as part of performance evaluation on AR.
@@
Here, we propose using tools from mechanistic interpretability to directly analyse the mechanisms used to solve synthetic tasks. We use causal interventions Geiger et al. [2024] on model internals to understand how these tasks are learned and implemented across a variety of architectures (section 4). This allows us to track the emergence (or lack thereof) of the correct association and retrieval mechanisms inside the model, beyond just observed task accuracy. Through comprehensive experiments on AR, we find that all SSMs except Based learn an inefficient direct-retrieval solution to AR, and that Mamba strongly relies on its short convolution component to perform AR.
@@
To deepen our findings, we introduce Associative Treecall (ATR), a novel synthetic retrieval task more similar to real-world natural language retrieval than AR (section 3). ATR uses a probabilistic context-free grammar (PCFG) to generate hierarchical data, on which we ask AR-like queries. Since keys and values need not be adjacent to each other, ATR requires a true non-positional retrieval mechanism, which may challenge architectures that are designed for AR. Interestingly, we observe the same mechanisms are implicated across architectures on ATR as on AR, indicating that association mechanisms are not task-dependent.
@@
Our results offer a framework for better understanding and evaluating synthetic task performance in terms of mechanistic interpretability. Mechanistic evaluations reveal fundamental differences between architectures beyond what we learn from behavioural performance, thus serving as a new tool for architecture analysis and design.
@@ 2 Related work
Associative Recall (AR) [^1] Also known as associative retrieval, associative memory, or induction. is a synthetic task that evaluates in-context retrieval for language model architectures, from early work on recurrent neural networks [Graves et al., 2014, Ba et al., 2016, Danihelka et al., 2016, Zhang and Zhou, 2017] to modern SSMs [Fu et al., 2023, Poli et al., 2023, Lutati et al., 2023, Jelassi et al., 2024, Arora et al., 2024a, b, Gu and Dao, 2024, Dao and Gu, 2024, Trockman et al., 2024, Liu et al., 2024a, Okpekpe and Orvieto, 2025, Li et al., 2025b, Wang et al., 2025]. An AR task consists of a sequence of key–value pairs followed by a single query key; the goal is to produce the corresponding value. For example,
@@
Here, the correct next token is 3, since it is the value associated with the key C in context. Despite being synthetic, AR has a direct analogue in natural language: induction, referring to in-context copying of sequences [Elhage et al., 2021, Olsson et al., 2022]. Arora et al. [2024a, b] show that architecture-level improvements on AR translate directly to natural-language induction.
@@
In order to measure the contribution of individual model components (neurons, layers, etc.) to output behaviour, we can apply causal interventions on neural network internals [Geiger et al., 2021, 2024]. Informally, the core idea is to overwrite an activation at a specific component using a counterfactual input. If this changes model behaviour, then that component is causally relevant to the mechanism underlying that behaviour.
@@
Some prior work in mechanistic interpretability has studied how some language models solve in-context retrieval tasks like induction and multiple choice question answering [Olsson et al., 2022, Lieberum et al., 2023, Brinkmann et al., 2024, Wiegreffe et al., 2025, Bick et al., 2025], as well as the training dynamics of Transformers on toy tasks using mechanistic metrics [Nanda et al., 2023, Reddy, 2024, Singh et al., 2024, Edelman et al., 2024, Tigges et al., 2024, Yin and Steinhardt, 2025]. Yet thus far, architectural comparisons on synthetic tasks have not made use of causal interventions.
3 Synthetic retrieval tasks
Induction, wherein key–value associations are stored in-context, is the memory-efficient mechanism implicated for retrieval tasks like AR in quadratic attention Transformers. Yet AR can also be solved through naïve positional association, and indeed SSMs theoretically learn a less efficient solution [Jelassi et al., 2024]. To elucidate this, we apply our mechanistic evaluation framework to compare architectures on two synthetic retrieval tasks: Associative Recall and Associative Treecall (ATR). Compared to AR, ATR is a novel language-like task with tree structure and more parameters for controlling task difficulty (section 3.2). Critically, ATR cannot be solved with naïve positional association, enabling us to explicitly test if models learn different mechanisms for association in the hierarchical setting. We build upon prior work on formal-language synthetic tasks [White and Cotterell, 2021, Valvoda et al., 2022, Hahn and Goyal, 2023, Strobl et al., 2024, Allen-Zhu and Li, 2024, Akyürek et al., 2024, Pandey, 2024, Lubana et al., 2024, inter alia].

### v3

@@ Abstract
State space models (SSMs) for language modelling promise an efficient and performant alternative to quadratic-attention Transformers, yet show variable performance on recalling basic information from the context. While performance on synthetic tasks like Associative Recall (AR) can point to this deficiency, behavioural metrics provide little information as to why—on a mechanistic level—certain architectures fail and others succeed. To address this, we conduct experiments on AR, and find that only Transformers and Based SSM models fully succeed at AR, with Mamba and DeltaNet close behind, while the other SSMs (H3, Hyena) fail. We then use causal interventions to explain why. We find that Transformers and Based learn to store key–value associations in-context using induction. By contrast, the SSMs seem to compute these associations only at the last state using a single layer. We further investigate the mechanism underlying the success of Mamba, and find novel evidence that Mamba does implement induction: not via the SSM, but instead via short convolutions. Further experiments on a new hierarchical retrieval task, Associative Treecall (ATR), show that all architectures learn the same mechanism as they did for AR. Furthermore, we show that Mamba can learn Attention-like induction on ATR when short convolutions are removed. These results reveal that architectures with similar accuracy may still have substantive differences, motivating the adoption of mechanistic evaluations.
@@ 1 Introduction
Transformers with quadratic attention remain the dominant architecture in language modelling despite numerous proposed efficient alternatives. Most notably, state space models (SSMs) achieve impressive perplexities and benchmark scores (e.g. Gu and Dao, 2024). Yet, SSMs exhibit deficiencies that benchmarks often fail to capture; for example, they struggle to perform retrieval, i.e. copying from the context Jelassi et al. (2024); Wen et al. (2024); Waleffe et al. (2024); Bick et al. (2025).
@@
Controlled synthetic tasks can make these limitations clear by isolating specific capabilities and enabling expressive experimentation at small scales across architectures. Particularly, much work has used the associative recall (AR) task as a testbed for studying in-context retrieval across architectures. In turn, AR has informed the design of novel LM architectures (e.g. Based; Arora et al., 2024b).
@@
However, in this prior work, performance on synthetic tasks is measured solely via behavioural metrics like task accuracy. This is a missed opportunity: an advantage of these tasks is that they are designed to isolate a specific behaviour that implicates a mechanistic solution. For example, LMs should solve AR by storing keyvalue associations in-context at the value. For Transformer-based LMs (Vaswani et al., 2017), the mechanism is generally thought to be the induction head. Olsson et al. (2022); Fu et al. (2023). We should therefore directly check whether each architecture learns an induction mechanism, as a way of checking that it has learned a robust solution to the task.
@@
Here, we propose using tools from mechanistic interpretability to directly analyse the mechanisms used to solve AR tasks. We use causal interventions Geiger et al. (2024) on model internals to understand how these tasks are learned and implemented across a variety of architectures (section˜3). This allows us to track the emergence (or lack thereof) of the correct association and retrieval mechanisms inside the model, which moves us beyond observed task accuracy and towards a detailed picture of the solutions different architectures learn.
Through comprehensive experiments on AR, we find that all SSMs except Based do not use induction in the traditional sense to implement AR (section˜4.2); instead, we find that the best-performing SSM, Mamba, relies heavily on the short convolution component in each layer to perform key-value association, and fails to learn AR at all without this (section˜4.3). Through layer-internal interventions, we discover that the short convolution component implements an induction-like associative mechanism in Mamba, with the SSM component only being used for retrieval (section˜4.4).
@@
To deepen our findings, we introduce Associative Treecall (ATR), a novel retrieval task more similar to real-world natural language retrieval than AR (section˜5). ATR uses a probabilistic context-free grammar (PCFG) to generate hierarchical data, on which we ask AR-like queries; the key difference from AR is that keys and their associated values need not be adjacent to each other. We find that the same mechanisms are implicated across architectures on ATR as on AR, confirming the generality of our findings (section˜5.2). However, unlike AR, Mamba does not need short convolutions to learn ATR; instead, mechanistic metrics reveal that Mamba falls back to an Attention-like two-layer induction mechanism when its short convolutions are removed (section˜5.3).
@@
We offer a framework for better understanding and evaluating task performance via mechanistic interpretability. We demonstrate multiple examples of mechanistic evaluations revealing fundamental differences between architectures that are unknown via behavioural performance, summarised in Figure˜1. We thus introduce mechanistic metrics as a new tool for architecture analysis.
Figure 1: Associative mechanisms in Attention, Based, and Mamba: Key results from applying mechanistic metrics to understand how Attention and SSMs solve AR and our new retrieval task, Associative Treecall (ATR). Attention and Based both implement induction, but with different architectural components (sections˜4.2 and 5.2). Mamba instead uses a single layer for both association and retrieval, but uses short convolution for association just like Based (sections˜4.3 and 4.4). When convolutions are removed, Mamba implements induction, but only on ATR (section˜5.3).
@@ 2 Related work
Our work relates to a variety of extant synthetic retrieval tasks. The foremost example, Associative Recall (AR), [^1] Also known as associative retrieval, associative memory, or induction. is a synthetic task that evaluates in-context retrieval for language model architectures. AR has been used extensively, from early work on recurrent neural networks (Graves et al., 2014; Ba et al., 2016; Danihelka et al., 2016; Zhang and Zhou, 2017) to modern SSMs (Fu et al., 2023; Poli et al., 2023; Lutati et al., 2023; Jelassi et al., 2024; Arora et al., 2024a; b; Gu and Dao, 2024; Dao and Gu, 2024; Trockman et al., 2024; Liu et al., 2024a; Okpekpe and Orvieto, 2025; Li et al., 2025b; Wang et al., 2025).
An AR task consists of a sequence of key–value pairs followed by a single query key; the goal is to produce the corresponding value for the given query. For example,
@@
Here, the correct next token is 3, since it is the value associated with the key C in context.
Despite being synthetic, AR has a direct analogue in natural language: induction, referring to in-context copying of sequences (Elhage et al., 2021; Olsson et al., 2022). Arora et al. (2024a; b) show that architecture-level improvements on AR translate directly to induction. Association is additionally widely studied in cognitive science as binding. Binding in neural networks has been examined by Greff et al. (2020); Kim and Schuster (2023); Feng and Steinhardt (2024); Prakash et al. (2024); Li et al. (2025a); Prakash et al. (2025).
@@
In order to measure the contribution of individual model components (neurons, layers, etc.) to output behaviour, we can apply causal interventions on neural network internals (Geiger et al., 2021; 2024). Informally, the core idea is to overwrite an activation at a specific component using a counterfactual input. If this changes model behaviour, then that component is causally relevant to the mechanism underlying that behaviour.
@@
Some prior work in mechanistic interpretability has studied how some language models solve in-context retrieval tasks like induction and multiple choice question answering (Olsson et al., 2022; Lieberum et al., 2023; Brinkmann et al., 2024; Wiegreffe et al., 2025; Bick et al., 2025), as well as the training dynamics of Transformers on toy tasks using mechanistic metrics (Nanda et al., 2023; Reddy, 2024; Singh et al., 2024; Edelman et al., 2024; Tigges et al., 2024; Yin and Steinhardt, 2025). Yet thus far, architectural comparisons on synthetic tasks have not made use of causal interventions.
3 Mechanistic metrics
Figure 2: An example interhcange intervention where we corrupt the key (A) and attempt to restore it by intervening at the value token in an intermediate representation. We evaluate the downstream effect on the next-token prediction at the query (which should ideally predict the answer).
Behavioural metrics provide little information as to why certain architectures succeed or fail on tasks of interest. Mechanistic metrics, which directly measure how information flows across model components and token positions, can tell us how AR and similar tasks are being solved by different architectures, and thus help us understand failures. We illustrate our approach in Figure˜2.
We use interchange interventions (Geiger et al., 2021; 2024) to understand and measure how solutions to AR are implemented across architectures. We introduce this operation and define the resulting metrics for our tasks below. Our implementation uses the pyvene library Wu et al. (2024).
Interchange intervention.
Consider a language model $p(\cdot)$ and some input $\mathbf{b}$. We select a component $f$ inside that model which computes some internal representation $f(\mathbf{b})$ during the LM’s forward pass. Now, consider a counterfactual input $\mathbf{s}$ which produces a counterfactual representation $f(\mathbf{s})$ when processed by $f$. We want to understand what about the output of $p$ is dependent on $f$. Therefore, we perform an intervention which replaces the output $f(\mathbf{b})$ with that of $f(\mathbf{s})$ during the computation of $p(\mathbf{b})$, with the change propagating downstream. The result is notated $p_{f\leftarrow f^{*}}(\mathbf{b},\mathbf{s})$.
Concrete setup for AR and ATR.
We take $\mathbf{o}$ to be a ground-truth document from our data distribution and $\mathbf{c}$ to be a version of that document with exactly one important token corrupted: the key (see Figure˜2). This corruption significantly reduces task accuracy by removing information that is necessary to answer the AR query.
For each architecture, we intervene at both the input and output each of the following model components $f$: each layer block, each sequence-mixer (e.g. Attention blocks in each Transformer layer), and each state-mixer (an MLP, except in Mamba, which lacks this component). [^2] We are not limited to analysing only these components, however; we do experiment with mixer-internal representations in our analysis. We measure to what extent the intervention can restore the likelihood of the correct answer to the query, i.e. we compare restored likelihood $p_{f\leftarrow f^{*}}(y_{\textrm{true}}\mid\mathbf{c},\mathbf{o})$ with original likelihood $p(y_{\textrm{true}}\mid\mathbf{o})$ and corrupted likelihood $p(y_{\textrm{true}}\mid\mathbf{c})$.
4 Understanding AR with mechanistic metrics
We now deploy our mechanistic metrics (section˜3) on AR. We follow the methodology outlined in section˜4.1 to create a variety of datasets and train models with various architectures and hyperparameter configurations. See appendix˜D for additional experiments not included here.
4.1 Methodology
Datasets.
We generate synthetic pretraining and evaluation datasets for AR. For each setting, the trainset has $100,032$ examples and the eval/dev sets have $320$ examples. In AR, we use disjoint key and value vocabularies. In each document, we separate the document from the query with a divider token, and provide only a single query. Further details are in appendix˜C.
Figure 3: Associative recall: Likelihood of correct answer without any interventions, after corrupting the key, and after restoring representations at the layer 1 block input with interchange intervention, on AR with vocabulary size $8192$ and key–value count of $32$. SSMs (except for Based) and Transformers learn different mechanisms.
Models.
We pretrain small models from scratch. We use the exact architecture implementations from the zoology [^3] https://github.com/HazyResearch/zoology library (Arora et al., 2024b) as well as the DeltaNet implementation from fla [^4] https://github.com/fla-org/flash-linear-attention Yang and Zhang (2024), except for behaviour-preserving modification of the LM backbone to enable interventions with pyvene [^5] https://github.com/stanfordnlp/pyvene (Wu et al., 2024) on various model-internal components. The LM backbone for all architectures is the same, with pre-norm blocks of alternating sequence mixers and MLPs (except for Mamba, which has no MLP) followed by LayerNorm at the end. We experiment with the following architectures: Attention (Vaswani et al., 2017), BaseConv Arora et al. (2024a), Based Arora et al. (2024b), DeltaNet Yang et al. (2024), H3 Fu et al. (2023), Hyena Poli et al. (2023), and Mamba Gu and Dao (2024); further details are given in appendix˜A.
Training.
We minimise next-token prediction cross-entropy loss, and mask the loss on all tokens except the query. We use the AdamW optimiser with $\beta=(0.9,0.999),\epsilon=10^{-8}$ and no weight decay. We warm up the learning rate for the first $10\%$ of training and then follow a cosine decay schedule to 0 for the remainder of training. We train for $16$ epochs with a batch size of $32$.
Each experiment trains $\approx 200$ models over all hyperparameters. Runtime varies from $0.5$ to $5$ hours, depending on hardware, task, and architecture. Overall, we used <10,000<10,000 GPU-hours in total, on a cluster with various NVIDIA machines (GPU memory 12.312.3G to 143.8143.8G).
Behavioural metrics.
We report behavioural metrics given the model’s predicted probabilities over the vocabulary $\hat{\mathbf{y}}\in\mathbb{R}^{\lvert\Sigma\rvert}$ and the index of the single true answer $i$. Our main metric is likelihood of the correct answer: $\hat{\mathbf{y}}_{i}$. We additionally report accuracy in appendices: $\mathds{1}[\mathrm{arg\,max}(\hat{\mathbf{y}})=i]$.
4.2 (Most) SSMs do not solve AR with induction
We first run experiments on a standard AR task and show that interchange interventions empirically confirm architectural differences in how AR is solved. We fix the total number of unique keys and values in the vocabulary to be $8192$, and present $32$ key–value pairs in context. As noted above, our trainset includes $100,032$ examples. We vary model dimensionality in $\{16,32,64,128,256\}$ and sweep LR in the range $3\cdot 10^{-5},3\cdot 10^{-2}]$ for each architecture.
Behavioural results.
The leftmost panel of Figure˜3 demonstrates that answer likelihood on AR cleanly separates Attention, which achieves $100\%$ likelihood and accuracy at $d\geq 32$, from nearly all SSMs. Based solves AR near-perfectly with roughly the same dimension-wise scaling curve as Attention, achieving a maximum likelihood of $98.9\%$ (accuracy of $99.06\%$). However, Mamba and DeltaNet are close behind and clearly better than other SSMs at AR, albeit achieving a less-than-perfect $90.3\%$ likelihood ($91.25\%$ accuracy) at $d=256$.
Mechanistic analysis.
After corrupting the key, we restore the original representation at the layer 1 block input at each of various positions (key, value, and query) and observe the resulting likelihood of the true answer. High restored likelihood at the value token causally indicates that the model is performing induction, whereas query indicates direct retrieval by the layer 0 block and key indicates the layer 1 block, respectively. We report results on the LR-tuned checkpoints of each architecture at each model dimension.
Our results in Figure˜3 cleanly separate Attention and Based, which only perform induction, from other SSMs, which either perform direct retrieval (Mamba, DeltaNet, Hyena) or fail to learn the task (H3, BaseConv). SSMs perform direct retrieval at varying layers: the best-performing SSMs almost entirely perform direct retrieval at layer 0 via the query token, with only one Mamba checkpoint ($d=32$) delays retrieval to layer 1. Jelassi et al. (2024) show that direct retrieval in SSMs has asymptotically worse capacity than the induction solution, and this is reflected in performance on AR.
4.3 Short convolutions enable AR in Mamba and Based
(a) Original and intervention restoration scores for Mamba when varying conv. kernel size; Restoring at the value is not shown since none of the models benefited from it.
(b) Accuracy on AR for Based when varying conv. kernel size or using implicit long conv.
Figure 4: Ablating short convolution in Mamba and Based: Accuracy and interchange intervention results when ablating parameters of the short convolution component in Mamba and Based. Conv. size less than 22, using long conv., or no conv., all lead to near-zero performance in both models.
We have observed that Attention, Based, and Mamba are the highest-performing architectures on AR. However, their underlying mechanisms differ: Attention and Based learn induction, a 2-layer mechanism which stores keyvalue associations at the value token as an intermediate step, whereas Mamba uses direct retrieval, a 1-layer mechanism which writes the association to the query token.
Importantly, Based and Mamba share a key architectural component: short convolutions. Based is a hybrid model with alternating short convolution and linear attention layers, while Mamba applies a short convolution to the layer input before each SSM block. We hypothesise that this architectural component is necessary [^6] Since Hyena also has a short convolution, this may not be sufficient for good performance on association. for learning AR when using a subquadratic sequence mixer. We conduct experiments on AR where we shorten the convolution kernel size in Mamba (from the default $d_{\mathrm{conv}}=4$ to $\{3,2,1\}$, and deleting it) and replace the Based short convolution with implicitly-parametrised long convolution (Poli et al., 2023).
Results.
We report results of our ablations in Figure˜4. On Mamba (Figure˜4(a)), we find a step change in task accuracy when increasing $d_{\mathrm{conv}}$ from $1$ to $2$, which introduces previous token information and thus enables AR. Without short convolution, Mamba fails to learn AR. Also, no Mamba checkpoints store the key at the value (i.e. induction), so direct retrieval persists. Finally, besides $d_{\mathrm{conv}}<2$ like Mamba, implicit long convolution in Based also significantly harms AR performance (Figure˜4(b)). We thus conclude that short convolutions are necessary for performing AR in Mamba and Based.
4.4 Actually, Mamba does do induction
Figure 5: Interventions on short convolutions in Mamba: Likelihood of the correct answer for all Mamba checkpoints (varying model dimension and LR) on AR; (from left to right) original performance, after corrupting the key, restoring at the layer 0 short conv output at the value, same but at the next key, and same but at the layer 1 value token. (Some runs failed due to high LR.)
We have established that (a) Attention and Based perform AR via induction, whereas Mamba and other SSMs use direct retrieval, and (b) Mamba and Based require short convolutions to succeed at AR. However, we have not elucidated what ‘direct retrieval’ is; its only apparent distinction from induction is that it is implemented using a single layer, but what is happening in that layer is unclear.
We have already observed in Based that the association step of AR can be implemented with a short convolution, and then retrieval can be handled by a component with a longer receptive field (linear attention). But section˜4.2 shows that this results in an induction mechanism in Based: short convolution, just like Attention, performs association by moving information about the key to its associated value. Might short convolution in Mamba serve the same role?
Methodology.
We use mechanistic metrics to analyse each layer’s short convolution components in all of our Mamba checkpoints (sweeping model dimension and LR). To enable Mamba-internal interventions, we use a native PyTorch implementation of Mamba and load in the weights from our hardware-optimised training runs; we confirm that unintervened performance is unchanged. As in section˜4.2, we corrupt the key, and attempt to restore key information by intervening at the hidden states outputted by the short convolution component in each layer. We report the likelihood of the correct answer after intervention, averaged over $64$ inputs.
Results.
Our results in Figure˜5 show that short convolution does move the key information; however, only a single checkpoint ($d=32$, $\mathrm{LR}=10^{-2}$) actually moves key information to the value (as expected in induction, cf. Attention), and that too in layer 1; instead, we observe in the remaining $9$ checkpoints with non-negligible performance, the information is moved to the next key by the layer 0 short convolution. While moving both the key and value to the following key is unusual, it still satisfies our two-stage definition of induction, and thus we claim that Mamba too performs induction.
5 Mechanistic findings on AR generalise to a new retrieval task
We now examine whether our results generalise to a novel retrieval task: Associative Treecall (ATR). ATR is more similar to real-world natural language retrieval than AR because it uses a probabilistic context-free grammar (PCFG) to generate hierarchical data, on which we ask AR-like queries. Since key and value need not be adjacent to each other, ATR requires a non-positional associative mechanism, which may challenge architectures that are designed for AR.
5.1 Associative Treecall (ATR)